Suppose a rocket is fired from the earth's surface of radius r, so that it just escapes from the gravitational influence of the earth with velocity v, then v is given by


1
Expert's answer
2021-01-24T15:47:28-0500

By the definition, the escape velocity is given by:


"KE=PE,""\dfrac{1}{2}mv^2=GM_Emr,""v_{esc}=\sqrt{\dfrac{2GM_E}{r}}"